The lifetime of fruit flies follows a normal distribution with mean of 70 days and standard deviation of 18 days. Top 20% are considered long-lived. At least how many days a fruit fly must survive in order for it to be considered long lived?
Answer #1

Given that,

mean = = 70

standard deviation = =18

Using standard normal table,

P(Z > z) = 20%

= 1 - P(Z < z) = 0.20

= P(Z < z ) = 1 - 0.20

= P(Z < z ) = 0.80

z = 0.84 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x= 0.84*18+70

x= 85.12